OpenStudy (anonymous):

**Calc 2 Help Needed** Solve the initial value problem below: (x^2 + 81)dy/dx = 3 y(9)=0

OpenStudy (amistre64):

calc2 has the easy separables

OpenStudy (anonymous):

I know this needs to be separated, but how?

OpenStudy (amistre64):

\[(x^2+81)\frac{dy}{dx}=3\] \[\frac{dx}{(x^2+81)}\left((x^2+81)\frac{dy}{dx}=3\right)\] \[dy=\frac{3}{(x^2+81)}dx\]

OpenStudy (amistre64):

then the left is a tan-1 type integration

OpenStudy (anonymous):

ok so dy integrated is y

OpenStudy (amistre64):

\[x=9tan(t)\] \[dx=9sec^2(t)dt\] \[dy=\frac{3}{((9tan(t))^2+81)}dx\] \[dy=\frac{3*9sec^2(t)}{(81tan^2(t)+81)}dt\] \[dy=\frac{27sec^2(t)}{81(tan^2(t)+1)}dt\] \[dy=\frac{1}{39}\frac{sec^2(t)}{sec^2(t)}dt\] then again, it might be pretty easy unless i messed it someplace

OpenStudy (amistre64):

1/3 , not 1/39

OpenStudy (anonymous):

hmmm. I'm still lost. haha

OpenStudy (anonymous):

where did you get the x = 9tanx?

OpenStudy (amistre64):

\[y=\frac13t\] \[x = 9tan(t)\] \[\frac x9 = tan(t)\] \[tan^{-1}(\frac x9) = t\] \[y=\frac13 tan^{-1}(\frac x9)+C\] maybe lol

OpenStudy (amistre64):

were did I get it? from a trig substitution technique

OpenStudy (amistre64):

i know if I had 81tan^2(t) + 81 I can manimpulate it into a 81sec^2(t) making the integration that much simpler

OpenStudy (amistre64):

i got an A in typing class, honest :)

OpenStudy (anonymous):

haha.. well I'm horrible at calculus..

OpenStudy (amistre64):

it just takes getting used to that you can actually adjust things in math. We grow up thinking this thin is static and unforgiving

OpenStudy (amistre64):

when I first learned this stuff im like; How do they expect us to know to do that? :)

OpenStudy (anonymous):

that's how I feel now. The class was relatively easy until 2 weeks ago.

OpenStudy (anonymous):

ok so doesn't there need to be a constant K on this problem? which I would plug back in with y(9) = 0?

OpenStudy (amistre64):

i use C as an "c"onstant

OpenStudy (amistre64):

we can call it M or W or \(\Large \phi \) if we want to

OpenStudy (anonymous):

is the final answer just an integer?

OpenStudy (amistre64):

by integer i take it yo mean some real number value; let see :) \[y=\frac13 tan^{-1}(\frac x9)+C\] \[0=\frac13 tan^{-1}(\frac 99)+C\] \[0=\frac13 tan^{-1}(1)+C\] \[0=\frac13 \frac{\pi}{4}+C\] \[0=\frac{\pi}{12}+C\] \[-\frac{\pi}{12}=C\] therefore \[y=\frac13 tan^{-1}(\frac x9)-\frac{\pi}{12}\]

OpenStudy (amistre64):

the final answer is an equation that will solve the initial condition

OpenStudy (anonymous):

makes sense! somewhat! haha. thanks for your help, you're a dang genius.

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

I have another one similiar if you want to help??

OpenStudy (amistre64):

maybe in a minute or so if noone else takes it; i gotta rest form this one ...

OpenStudy (anonymous):

ok thanks!